OpenStudy (anonymous):

The speed of two pulleys varies inversely as their diameters. A pulley with a diameter of 32 centimeters turns at 600 revolutions per minute. What is the speed of a pulley with a diameter of 24 centimeters? The speed of the pulley is ___???___ revolutions per minute.

OpenStudy (lgbasallote):

inversely proportional so \[y = \frac kx\] let y = speed of pulley and x is diameter when y = 600, x = 32 \[\implies 600 = \frac{k}{32}\] \[\implies 600 \times 32 = k\] \[\implies 19, 200 = k\] so the qquestion is..what is y when x = 24 \[\implies y = \frac kx\] \[\implies y= \frac{19,200}{24}\] \[\implies y = 800\] does that help?

OpenStudy (anonymous):

yes yes yes !!!!!

OpenStudy (lgbasallote):

wonderful